OpenStudy (anonymous):

Which statement is an example of the Addition Property of Equality? (1 point) If p = q then p • s = q • s . If p = q then p + s = q + s . If p = q then p - s = q - s . p = q

OpenStudy (swissgirl):

Well the addition property of equality: If a = b then a + c = b + c.

OpenStudy (anonymous):

can u change that to p, and q and s? ;)

OpenStudy (swissgirl):

Yes u can

OpenStudy (swissgirl):

These are just dummy variables

OpenStudy (anonymous):

can u help me do it?

OpenStudy (swissgirl):

All its saying is that if two variables equal each other then by adding the same number to both it will still be equal to each other So for instance if m=n then even once adding 3 to both besides it is still equal m+3=n+3

OpenStudy (anonymous):

so the second one?

OpenStudy (swissgirl):

Ya

OpenStudy (anonymous):

thank u!!